OpenStudy (anonymous):

In a certain year the amount A of garbage in pounds produced after t days by an average household is given by A= 5.5t. Graph the equation t>0. How many days did it take for the average household to produce 200 pounds of garbage? Please help. So confused.

hartnn (hartnn):

in the equation A=5.5t A is given as 200,u have to find t so,200=5.5t can u solve further to get t?

OpenStudy (anonymous):

does this mean that I subtract 5.5t from both sides

hartnn (hartnn):

to isolate t,u divide 5.5 on both sides,its just like this a=bc gives c=a/b

OpenStudy (anonymous):

so this means...If I divide both sides 5.5t/5.5 = 1t

OpenStudy (anonymous):

and 200/ 5.5= 36.36

hartnn (hartnn):

exactly :) so u got t as36.36 days ..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

to check my answer I would fill in the equation..200= 5.5 x 36.36

hartnn (hartnn):

yes,right and you would get both sides equal,indicating your answer is correct....

OpenStudy (anonymous):

I appreciate your help

hartnn (hartnn):

anytime :)

OpenStudy (anonymous):

: )